# Quantifying the Uniqueness of Donald Trump in Presidential Discourse

### v1

@@ Appendix B Additional LM-based Uniqueness Results
    Figure 10: Heatmaps of Pearson Correlation Coefficients between our uniqueness metric and existing readability scores (FKGL, Gunning Fog, Dale Chall, and ARI). There is little to no correlation between our score and readability; in contrast, there is strong correlation between the different readability indexes.
    B.1 BPC
    Recall that we employ bits-per-character (BPC), also known as bits-per-byte, as a proxy for “predictability” of a sentence. Lower BPC values correspond to higher predictability. We use BPC instead of perplexity or loss directly to account for variation in tokenization techniques. BPC is the sum of cross-entropy losses of each token in a sentence, divided by the number of characters (bytes) in the sentences. Fig. 9 shows average sentence BPC per candidate (Fig. 8(a)) and aggregated by party over sentence length (Fig. 8(b)). While in campaigns, Trump is the most unpredictable among those candidates, he is ranks closer to the middle for predictability in debates and SOTU.
    B.2 Correlation with Readability Metrics
    We calculate the Pearson correlation coefficients for our uniqueness metric with existing readability scores. Our metric has little to no correlation with readability, while different readability indexes show strong correlation with each other (Fig. 10). Fig. 10 also confirms that our metric is dintinct from sentence length.
    B.3 Uniqueness scores over time
@@ B.4 Top %-ile uniqueness
    B.5 LM-based uniqueness results with unmasked model
    Appendix C Additional Divisive Word Lexicon Results
    C.1 Divisive Word Lexicon
    Table 5 contains the 178 words in our proposed divisive word lexicon.
@@ C.1 Divisive Word Lexicon
    Table 5: The 178 words in our divisive word lexicon, which are verified as being divisive by the majority of 4 annotators. Our original ten seed words are denoted with a $*{}^{*}$.
    C.2 Usage over time
@@
    C.3 Usage by candidate
    C.4 Utterances Using Divisive Words
@@
    2020
    Joe Biden: I’m talking about the Biden plan. . .
@@
    2020
    Joe Biden: Russia is paying you a lot.
@@
    2020
@@
    2020
    Donald Trump: I mean, they can say anything.
@@
    2016
    Donald Trump: Well, all of these bad leaders from ISIS are leaving Mosul.
@@
    2016
    Donald Trump: But the leaders that we wanted to get are all gone because they’re smart.
@@
    1976
    Gerald Ford: On the other hand, when you have a bill of that magnitude, with those many provisions, a President has to sit and decide if there is more good than bad.
@@
    1988
    George Bush: In terms of negative campaigning, you know, I don’t want to sound like a kid in the schoolyard: he started it.
@@
    2022
@@
    2020
    Donald Trump: The terrorist responsible for killing Sergeant Hake was Qasem Soleimani, who provided the deadly roadside bomb that took Chris’s life.
@@
    2019
    Donald Trump: On Friday, it was announced that we added another 304,000 jobs last month alone, almost double the number expected.
@@
    2018
    Donald Trump: In April, this will be the last time you will ever file under the old and very broken system, and millions of Americans will have more take-home pay starting next month—a lot more.
@@
    2017
@@
    2016
    Barack Obama: But after years now of record corporate profits, working families won’t get more opportunity or bigger paychecks just by letting big banks or big oil or hedge funds make their own rules at everybody else’s expense.
@@
    2012
@@
    1991
    George Bush: Last year, our friends and allies provided the bulk of the economic costs of Desert Shield.
@@
    1982
    Ronald Reagan: Contrary to some of the wild charges you may have heard, this administration has not and will not turn its back on America’s elderly or America’s poor.
@@
    2020
    Joe Biden: The last thing you need is a President who ignores you, looks down at you, who just doesn’t understand you.
@@
    2020
    Joe Biden: I’m so grateful to have earned the UA’s endorsement — and to have 355,000 proud plumbers, pipefitters, and more behind me.
@@
    2020
    Donald Trump: This is the most important election in the history of our country.
@@
    2016
@@
    2016
@@
    2016
@@
    2016
    Hillary Clinton: We should honor the men and women in uniform who fight for our country.
@@
    2008
    John McCain: In a time of trouble and danger for our country, who will put our country first?
@@
    2008
    John McCain: In his three short years in the Senate, he has requested nearly a billion dollars in pork projects for his state - a million dollars for every day he’s been in office.
@@
    Appendix D Additional Opponent Mention Results
    D.1 Rate of opponent mentions
    D.2 Fightin’ Words overlap

### v2

@@ D.2 Fightin’ Words overlap
Figure 11: Heatmaps of Spearman Correlation Coefficients between our uniqueness metric and existing readability scores (FKGL, Gunning Fog, Dale Chall, and ARI). There is little to no correlation between our score and readability; in contrast, there is strong correlation between the different readability indexes.
@@
(a) Gemma 2B
(b) Phi-1.5
Figure 18: Validation of the LLM-based uniqueness metric with alternative LLMs. The resulting trends are consistent with original findings that Trump’s speech is more unique than other Democratic and Republican presidential candidates.
stupid,
dishonest,
unamerican,
idiot,
deplorable,
pathetic,
immoral,
disgrace,
incompetent,
foolish,
@@
Table 6: The 178 words in our divisive word lexicon, which are verified as being divisive by the majority of 4 annotators. Our original ten seed words are denoted with a .
@@
2020 Joe Biden: I’m talking about the Biden plan. . .
@@
2020 Joe Biden: Russia is paying you a lot.
@@
2020 Donald Trump: We have to go back to the core values of this country.
@@
2020 Donald Trump: I mean, they can say anything.
@@
2016 Donald Trump: Well, all of these bad leaders from ISIS are leaving Mosul.
@@
2016 Donald Trump: But the leaders that we wanted to get are all gone because they’re smart.
@@
2016 Hillary Clinton: And we should demand that Donald release all of his tax returns so that people can see what are the entanglements and the financial relationships that he has…
Other: We’re going to get to that later.
Other: Secretary Clinton, you’re out of time.
Hillary Clinton: … with the Russians and other foreign powers.
Other: Mr. Trump?
Donald Trump: Well, I think I should respond, because so ridiculous.
1976 Gerald Ford: On the other hand, when you have a bill of that magnitude, with those many provisions, a President has to sit and decide if there is more good than bad.
@@
1988 George Bush: In terms of negative campaigning, you know, I don’t want to sound like a kid in the schoolyard: he started it.
@@
2022 Joe Biden: Putin is now isolated from the world more than he has ever been.
Joe Biden: Together, along with our allies, we are right now enforcing powerful economic sanctions.
@@
2020 Donald Trump: The terrorist responsible for killing Sergeant Hake was Qasem Soleimani, who provided the deadly roadside bomb that took Chris’s life.
@@
2019 Donald Trump: On Friday, it was announced that we added another 304,000 jobs last month alone, almost double the number expected.
@@
2018 Donald Trump: I am asking Congress to address the fundamental flaws in the terrible Iran nuclear deal.
Donald Trump: My administration has also imposed tough sanctions on the communist and socialist dictatorships in Cuba and Venezuela.
Donald Trump: But no regime has oppressed its own citizens more totally or brutally than the cruel dictatorship in North Korea.
2018 Donald Trump: In April, this will be the last time you will ever file under the old and very broken system, and millions of Americans will have more take-home pay starting next month—a lot more.
@@
2017 Donald Trump: We cannot allow a beachhead of terrorism to form inside America.
Donald Trump: We cannot allow our Nation to become a sanctuary for extremists.
Donald Trump: That is why my administration has been working on improved vetting procedures, and we will shortly take new steps to keep our Nation safe and to keep those out who will do us harm.
@@
2016 Barack Obama: But after years now of record corporate profits, working families won’t get more opportunity or bigger paychecks just by letting big banks or big oil or hedge funds make their own rules at everybody else’s expense.
@@
2012 Barack Obama: In 2008, the house of cards collapsed.
Barack Obama: We learned that mortgages had been sold to people who couldn’t afford or understand them.
@@
1991 George Bush: Last year, our friends and allies provided the bulk of the economic costs of Desert Shield.
@@
1982 Ronald Reagan: Contrary to some of the wild charges you may have heard, this administration has not and will not turn its back on America’s elderly or America’s poor.
@@
2020 Joe Biden: The last thing you need is a President who ignores you, looks down at you, who just doesn’t understand you.
@@
2020 Joe Biden: I’m so grateful to have earned the UA’s endorsement — and to have 355,000 proud plumbers, pipefitters, and more behind me.
@@
2020 Donald Trump: This is the most important election in the history of our country.
@@
2016 Donald Trump: Now, Bernie Sanders should be angry right?
Donald Trump: Shouldn’t he be angry?
Donald Trump: Now, I’ll tell you what.
Donald Trump: The system is rigged.
@@
2016 Donald Trump: I want the entire corrupt Washington establishment to hear and to heed the words we all will be saying right now.
Donald Trump: When we win on November 8th, we are going to Washington, D.C. and we are going to drain the swamp.
Donald Trump: Gonna drain the swamp.
@@
2016 Donald Trump: Our trade deals, we lose \$800 billion a year on trade.
@@
2016 Donald Trump: We will terminate NAFTA and get a much better deal for our workers if we can’t renegotiate it properly.
Donald Trump: We’re gonna get a better deal for our workers and for our companies.
Donald Trump: Because we cannot continue to be the people led by stupid people.
2016 Hillary Clinton: We should honor the men and women in uniform who fight for our country.
@@
2016 Hillary Clinton: But this is not new.
Hillary Clinton: I know I’m reaching out to Republicans and Independents as well as Democrats because I want to be the president for all Americans.
Hillary Clinton: And, when you think about it, what he said at the convention, ’I alone can fix it,’ runs counter to who we are as Americans.
Hillary Clinton: We work together.
Hillary Clinton: So there are many reasons why I think it is fair to conclude that Donald Trump is unqualified and unfit to be president.
2008 John McCain: In a time of trouble and danger for our country, who will put our country first?
@@
2008 John McCain: In his three short years in the Senate, he has requested nearly a billion dollars in pork projects for his state - a million dollars for every day he’s been in office.
@@
(a) Sentence uniqueness across divisive language usage, for each type of speech.
(b) Sentence uniqueness across opponent mentions, in debates.
(c) Divisive language frequency across opponent mentions, in debates
Figure 23: Pairwise analysis of our three proposed metrics, aggregated globally across all speakers. Fig. 23(a) shows that in debates and campaign speeches, sentences that use divisive language tend to be more unique as well. For sentences containing opponent mentions in debates, Fig. 23(b) shows that such utterances tend to be more distinctive and Fig. 23(c) shows that they tend to have higher divisive word usage as well.